# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ II.3 Open heavy-flavor data
Open heavy-flavor production in $h+A$ collisions offers complementary insights into nuclear matter effects relative to quarkonium production. At LO, heavy quarks are predominantly produced through $2\to 2$ scatterings. Following production, they undergo hadronization via fragmentation into various charged and neutral states, including $D^{0}$, $D^{+}$, and $D^{-}$ mesons for charm quarks, and $B^{0}$, $B^{+}$, $B^{-}$ mesons for bottom quarks. The corresponding nuclear modification factors have been investigated both at fixed-target and collider facilities.
The first experiments were with pion beams. WA82 at CERN [113] and E769 at Fermilab [114] investigated the $A$ dependence of $D^{\pm}$, $D^{0}$ and $\bar{D}^{0}$ mesons. The resulting cross sections were assumed to follow the usual $A^{\alpha}$ dependence. In both experiments, the extracted values of $\alpha$ was found to be compatible with unity. Moreover, no statistically significant dependence on $p_{T}$ or $x_{F}$ was observed. A subsequent analysis of $D^{*}$ mesons by E769 [115] corroborated these findings. An absence of nuclear effects was also reported [116] by the WA92 experiment at CERN for $D^{0}$, $D^{+}$, and $D_{s}^{+}$ mesons and their charge conjugates. Similarly, a lack of nuclear dependence in the central rapidity region (near $x_{F}=0$) was reported by the E789 collaboration [117], using an 800 GeV proton beam. More recently, these results were confirmed and their precision improved by the HERA-B collaboration for $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $D^{*+}$ mesons in $p+A$ collisions at 920 GeV [118].
@@
While no significant $A$ dependence was observed for $D$ mesons, a marked asymmetry was observed between $D^{+}$ and $D^{-}$ meson distributions as a function of $x_{F}$ and $p_{T}$. With a $\pi^{-}$ beam, the $D^{-}$, which shares a valence $d$ quark with the beam, is designated as leading charm while the $D^{+}$, which does not, is considered non leading. The E791 collaboration [120] compared their $D$ meson asymmetries with those from E769 [114] and WA82 [113] and found that, even though the energies were different, the asymmetries were similar. More recently, SMOG@LHCb has demonstrated a similar asymmetry in $p+{\rm Ne}$ interactions [121]. These data cannot be easily explained within convention production in pQCD.
Open charm data were also collected by the LHCb collaboration in fixed-target mode [86], using the LHC proton beam at 2.5 TeV and a gaseous neon target. A comparison between $J/\psi$ and $D^{0}$ production shows a cross section ratio that more than doubles for 1 $\lesssim p_{T}\lesssim$ 4 GeV, while it remains nearly constant as a function of the rapidity.
Additional measurements have been made at the LHC by the ALICE [122, 123] and LHCb collaborations [124, 125] at $\sqrt{s_{NN}}=5.02$ TeV, with additional $D^{0}$ and $B$ meson data from LHCb at $\sqrt{s_{NN}}=8.16$ TeV [126, 127]. The LHCb measurements reveal distinct rapidity-dependent effects. In $p+\mathrm{Pb}$ collisions, $R_{pA}^{D}$ exhibits a weak $p_{T}$ dependence at backward rapidity, while strong suppression is observed at forward rapidity with low $p_{T}$. Similarly, $B^{+}$ meson production at $\sqrt{s_{NN}}=8.16$ TeV shows a more pronounced suppression at forward rapidity compared to backward rapidity. Low $p_{T}$ suppression is observed in the forward region while there is no significant suppression at backward rapidity.
In addition, LHCb collaboration has reported the first observation of strangeness enhancement in charm hadronization in $p+\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=8.16$ TeV [128]. This effect was studied through the ratio of $D_{s}^{+}$ to $D^{+}$ mesons as a function of charged-particle multiplicity, measured in both the backward ($-5.0<y<-2.5$) and forward ($4.01.5<y<$ 4.0) rapidity regions. A significant enhancement of $D_{s}^{+}$ over $D^{+}$ production is observed in high-multiplicity events, particularly at low transverse momentum and in the backward rapidity region.
Recently, the ratio of charm baryon $\Lambda_{c}^{+}$ to charm meson $D^{0}$ production in proton-lead collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ was published by the CMS collaboration as a function of $p_{T}$ in the rapidity range $-1.46<y<0.54$. A nuclear dependence of charm baryons relative to charm mesons is observed [129, 130]. A rapidity-dependent analysis could further elucidate the underlying interaction dynamics.
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Process Target ($A$) Dependence Ref.
@@ II.4 Ultraperipheral collisions in ultrarelativistic heavy-ion collisions
Ultraperipheral heavy-ion collisions (UPCs) provide a unique experimental tool to generate large cross sections of high-energy photon-nucleus collisions and characterize the nature of nuclear matter at low $x$ [133]. As shown in Fig. 1, UPCs are characterized by an impact parameter that exceeds the sum of the two colliding nuclear radii. These interactions are dominated by scatterings of quasi-real photons from the strong electromagnetic fields accompanying ultrarelativistic heavy nuclei. [^3] In such collisions, the photon flux from each nucleus is proportional to the square of the nuclear charge, $Z^{2}$. The study of heavy-quark production in UPCs, in particular, relies on perturbative calculations to model the partonic scattering.
Measurements of coherent ${J}/\psi$ photoproduction in UPCs by STAR [134] at RHIC and ALICE [135, 136, 137] and CMS [138] at the LHC could constrain the gluon nPDF in nuclei for $Q^{2}\sim m^{2}_{c\bar{c}}$ as low as $x\approx 10^{-5}$, confirming strong suppression of low $x$ gluons [139, 140]. More recently, the first measurement of incoherent ${J}/\psi$ production in UPCs [141] has provided insights into event-by-event gluon density fluctuations. CMS results show significant nuclear suppression, comparable to coherent photoproduction, for $x<10^{-4}$. The suppression decreases with increasing $x$.
@@
Experiment System $\sqrt{s_{NN}}$ (GeV) Process Dependence Reference
@@ II.5 Discussion
We have presented a non-exhaustive overview of the behavior of nuclear modification factors for several hard processes. With the exception of some DY data, suppression or enhancement patterns are clearly visible in all the energy domain covered. However, the extent of the observed modification depends on the specific process and observable under consideration.
Here we highlight the key observations and considerations that illustrate the limitations of nPDF formalism—which models the modification of parton distributions inside nuclei relative to free nucleons—in fully describing the $h+A$ data in Tables 1, 2, 3, and 4 as a function of rapidity ($y$) and transverse momentum ($p_{T}$). Even in the case of structure functions in DIS, it remains unclear how much of the observed behavior arises from leading-twist dynamics as opposed to nuclear-enhanced higher-twist effects [148, 149, 150]. The nPDF formalism will be discussed in more detail in Section III.1. The nonexhaustive list below identifies some strong indications that going beyond the nPDF framework is necessary to fully interpret the data.
  1. (i)
    In the $2\to 1$ kinematics regime ($p_{T}\lesssim M$), relevant for processes like $gg\to J/\psi$ or $q\bar{q}\to\gamma^{\star}$ (Drell-Yan), the dimuon $p_{T}$ is negligible in terms of $x_{2}$ and $Q^{2}$. As a result, the nuclear modification factor in hadron-nucleus collisions simplifies to
@@
    Figure 2: The $p_{T}$ dependence of the Drell-Yan ($8.54<M<$ 8.5 GeV, blue) and $J/\psi$ ($0.3\leq x_{\text{F}}\leq 0.93$, red) nuclear modification factorin $p+A$ collisions at $\sqrt{s_{NN}}=38.7$ GeV, as measured by the E866 experiment [40, 80].
  2. (ii)
@@
  4. (iv)
    At LHCb energies, $R_{pA}^{J/\psi}\sim 0.6$ at forward rapidity ($42.5<y<$ 4) [100, 103], stronger than predicted by nPDF shadowing. This suppression is consistent with observations from the ALICE collaboration [94, 97], which reported a similar trend in $J/\psi$ suppression at forward rapidity in $p+\mathrm{Pb}$ collisions at the same beam energy. In addition, the observed hierarchy in nuclear modification factors, $R_{pA}^{\psi{\rm(2S)}}<R_{pA}^{J/\psi}<R_{pA}^{\Upsilon(1S)}$ [110, 112], suggests the presence of nuclear effects beyond nPDFs alone.
  5. (v)
@@
    Figure 4: Nuclear dependence of (top) Drell-Yan and (bottom) $J/\psi$ production as a function of $x_{2}$. The Drell-Yan data, $R_{pA}^{\rm DY}$, are from E772 [39], E866 [40], and E906 [44]. The $J/\psi$ data, presented as the exponent $\alpha$, are from NA3 [73], E866 [80], PHENIX [153], ALICE [154], and LHCb [155]. Adapted from [42].
All these observations, across a range of $x_{F}$, $x_{2}$ and $p_{T}$ domains, suggest the need to go beyond the conventional nPDF framework and demonstrate an energy dependence of nuclear effects. We attempt to identify the physics mechanisms that could provide a complete and universal description of these data. The following sections introduce the cold nuclear matter effects that may underlie the observed experimental trends and aim to address the three key questions outlined in Section I. We also propose a set of observables designed to disentangle contributions to the observed behavior from individual nuclear effects, thereby facilitating a more precise characterization of the nuclear modifications.
@@ III Disentangling nPDFs from saturation
If this suppression is genuinely nonperturbative, one must then rely on phenomenological model and possibly lattice methods, which currently face severe limitations. One such model is the Glauber-Gribov model [159, 160, 161, 162, 163] of (virtual) photon-nucleus total cross sections (structure functions) where the exchanged degrees of freedom are non-perturbative color singlet states called “pomerons” [164, 165] (or soft pomerons, to be distinguished from hard or BFKL pomerons [166]) which carry the quantum numbers of vacuum. In this approach, suppression of the nuclear cross section is due to the destructive interference between the phases of multiple pomeron exchanges. The nuclear parton distributions are thus determined at an initial scale $Q_{0}^{2}$ and then evolved in $Q^{2}$ via the DGLAP equations [167, 168]; this procedure is known as leading-twist shadowing. The evolution of the parton distribution functions is identical for nucleons and nuclei, with all nuclear effects (shadowing) encoded in the initial distributions. These nuclear PDFs, assumed to be universal, are then used within the collinear factorization framework to compute particle production cross sections in nuclear collisions.
@@
On the other hand, if weakly coupled QCD dynamics underlie the observed suppression, it should be calculable directly in terms of quark and gluon degrees of freedom. One such mechanism is gluon saturation, formulated as the Color Glass Condensate (CGC) effective field theory of small $x$ [169, 170, 171]. Such a saturated state of gluons, characterized by a relatively hard scale, the so-called saturation scale $Q_{s}\gg\Lambda_{\rm QCD}$, can be reached in a large nucleus already at moderately small values of $x$ ($<4-5\times 10^{-2}$) without significant evolution in $x$ whereas, in the proton, saturation is expected to be reached at much smaller values of $x$. In this regime of high gluon densities, $Q^{2}\lesssim Q_{s}^{2}$, nonlinear gluon recombination effects play a crucial role in taming the rapid growth of the gluon distribution at low $x$, leading to saturation of the distribution, as required by unitarity.
@@
The saturation scale, which delineates the transition between the dilute and dense regimes of QCD matter, depends on the atomic number $A$, the gluon momentum fraction $x$ (which coincides with Bjorken $x$ in structure functions but not necessarily in particle production), and the impact parameter of the nucleus or proton. Importantly, as an effective theory of QCD at small $x$, the CGC framework applies to a broad class of observables beyond structure functions [172, 173].
@@
Note that CGC dynamics can lead to both higher twist and leading twist suppression of nuclear cross sections. In the kinematic regime where small-$x$ evolution is significant but the momentum or virtuality remains well above the saturation scale ($Q^{2}\gg Q_{s}^{2}$), the cross section can be expanded in powers of $1/Q^{2}$, with the leading contribution corresponding to leading-twist dynamics. Although the system is dilute in this regime, leading-twist shadowing arises due to the presence of a saturation boundary separating the dense and dilute regimes. The extent of the suppression is sensitive to the dynamics near the saturation scale, which can be predicted within the CGC effective theory. Lastly, there has been significant progress toward calculating next-to-leading order (NLO) corrections for various observables using the CGC formalism which is crucial for precise predictions [174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207].

### v2

@@ II.3 Open heavy-flavor data
Open heavy-flavor production in $h+A$ collisions offers complementary insights into CNM effects relative to quarkonium production. At LO, heavy quarks are predominantly produced through $2\to 2$ scatterings. Following production, they undergo hadronization via fragmentation into various charged and neutral states, including $D^{0}$, $D^{\pm}$, and $D_{s}$ mesons for charm quarks, and $B^{0}$, $B^{\pm}$, $B_{c}$ and $B_{s}$ mesons for bottom quarks. The corresponding nuclear modification factors have been investigated both at fixed-target and collider facilities.
The first experiments were with pion beams. WA82 at CERN [113] and E769 at Fermilab [114] investigated the $A$ dependence of $D^{\pm}$, $D^{0}$ and $\bar{D}^{0}$ mesons. The resulting cross sections were assumed to follow the usual $A^{\alpha}$ dependence. In both experiments, the extracted values of $\alpha$ were found to be compatible with unity. Moreover, no statistically significant dependence on $p_{T}$ or $x_{F}$ was observed. A subsequent analysis of $D^{*}$ mesons by E769 [115] corroborated these findings. An absence of nuclear effects was also reported [116] by the WA92 experiment at CERN for $D^{0}$, $D^{+}$, and $D_{s}^{+}$ mesons and their charge conjugates. Similarly, no nuclear dependence was observed in the central rapidity region (near $x_{F}=0$) by the E789 collaboration [117], using an 800 GeV proton beam. More recently, these results were confirmed and their precision improved by the HERA-B collaboration for $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $D^{*+}$ mesons in $p+A$ collisions at 920 GeV [118].
@@
While no significant $A$ dependence was observed for $D$ mesons, a marked asymmetry was observed between $D^{+}$ and $D^{-}$ meson distributions as a function of $x_{F}$ and $p_{T}$. With a $\pi^{-}$ beam, the $D^{-}$, which shares a valence $d$ quark with the beam, is designated as leading charm while the $D^{+}$, which does not, is considered nonleading. The E791 collaboration [120] compared their $D$ meson asymmetries with those from E769 [114] and WA82 [113] and found that, even though the energies were different, the asymmetries were similar. More recently, SMOG@LHCb has demonstrated a similar asymmetry in $p+{\rm Ne}$ interactions [121]. These data cannot be easily explained within convention production in pQCD but could be interpreted as being due to effects such as intrinsic charm or color connections.
Open charm data were also collected by the LHCb collaboration in fixed-target mode [86], using the LHC proton beam at 2.5 TeV and a gaseous neon target. A comparison between $J/\psi$ and $D^{0}$ production shows a cross section ratio that more than doubles for 1 $\lesssim p_{T}\lesssim$ 4 GeV, while it remains nearly constant as a function of $y$.
Measurements in collider mode have also been made at the LHC by the ALICE [122, 123] and LHCb collaborations [124, 125] at $\sqrt{s_{NN}}=5.02$ TeV, with additional $D^{0}$ and $B$ meson data from LHCb at $\sqrt{s_{NN}}=8.16$ TeV [126, 127]. The LHCb measurements reveal distinct rapidity-dependent effects. In $p+\mathrm{Pb}$ collisions, $R_{pA}^{D}$ exhibits a weak $p_{T}$ dependence at backward rapidity, while strong suppression is observed at forward rapidity with low $p_{T}$. Similarly, $B^{+}$ meson production at $\sqrt{s_{NN}}=8.16$ TeV shows a more pronounced suppression at forward rapidity compared to backward rapidity. Low $p_{T}$ suppression is observed in the forward region while there is no significant suppression at backward rapidity.
In addition, the LHCb collaboration has reported the first observation of strangeness enhancement in charm hadronization in $p+\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=8.16$ TeV [128] by measuring the $D_{s}^{+}$ to $D^{+}$ ratio as a function of charged-particle multiplicity in both the backward ($-5.0<y<-2.5$) and forward ($4.01.5<y<$ 4.0) rapidity regions. A significant enhancement of $D_{s}^{+}$ over $D^{+}$ production is observed in high-multiplicity events, particularly at low $p_{T}$ and backward rapidity.
Recently, the ratio of the $\Lambda_{c}^{+}$ charm baryon relative to $D^{0}$ production in $p+{\rm Pb}$ collisions at $\sqrt{s}=5.02$ TeV was published by the CMS collaboration as a function of $p_{T}$ for $-1.46<y<0.54$. An $A$ dependence of charm baryon relative to charm meson production is observed [129, 130]. A rapidity-dependent analysis could further elucidate the underlying interaction dynamics.
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Final State Target ($A$) Observable Ref.
@@ II.4 Ultraperipheral collisions in ultrarelativistic heavy-ion collisions
Ultraperipheral heavy-ion collisions (UPCs) provide a unique experimental tool to generate large fluxes of high-energy photons for $\gamma+A$ collisions to provide further characterization the nature of nuclear matter at low $x$ [133]. As shown in Fig. 1, UPCs are characterized by an impact parameter that exceeds the sum of the two colliding nuclear radii. These interactions are dominated by scatterings of quasi-real photons from the strong electromagnetic fields accompanying ultrarelativistic heavy nuclei. [^3] In such collisions, the photon flux from each nucleus is proportional to the square of the nuclear charge, $Z^{2}$. The study of heavy-quark production in UPCs, in particular, relies on perturbative calculations to model the partonic scattering.
Measurements of coherent ${J}/\psi$ photoproduction in UPCs by STAR [134] at RHIC and ALICE [135, 136, 137] and CMS [138] at the LHC could constrain the gluon nPDF in nuclei for $Q^{2}\sim M^{2}_{c\overline{c}}$ as low as $x\approx 10^{-5}$, confirming strong suppression of low $x$ gluons [139, 140]. More recently, the first measurement of incoherent ${J}/\psi$ production in UPCs [141] provided insights into event-by-event gluon density fluctuations. CMS results show significant nuclear suppression, comparable to coherent photoproduction, for $x<10^{-4}$. The suppression decreases with increasing $x$.
@@
Experiment System $\sqrt{s_{NN}}$ (GeV) Final State Observable Reference
@@ II.5 Discussion
We have presented a non-exhaustive overview of the behavior of nuclear modification factors for several hard processes. With the exception of some Drell-Yan data, suppression or enhancement patterns are clearly visible in all the energy domains covered. However, the extent of the observed modification depends on the specific process and observable under consideration.
Here we highlight the key observations and considerations that illustrate the limitations of the nPDF formalism, which models the modification of parton distributions inside nuclei relative to free nucleons, in fully describing the $h+A$ data in Tables 1, 2, 3, and 4 as a function of $y$ and $p_{T}$. Even in the case of structure functions measured in DIS, it remains unclear how much of the observed behavior arises from leading-twist dynamics as opposed to nuclear-enhanced higher-twist effects [148, 149, 150]. The nPDF formalism will be discussed in more detail in Section III.1. The nonexhaustive list below identifies some strong indications that going beyond the nPDF framework is necessary to fully interpret the data.
  • In the $2\to 1$ LO kinematics regime ($p_{T}\lesssim M$), relevant for $gg\to J/\psi$ or $q\bar{q}\to\gamma^{\star}$ (Drell-Yan), the final-state $p_{T}$ is negligible for the determination of $x_{2}$ and $Q^{2}$. As a result, the nuclear modification factor in hadron-nucleus collisions simplifies to
@@
    Figure 2: The $p_{T}$ dependence of the Drell-Yan ($8.54<M<$ 8.5 GeV, blue) and $J/\psi$ ($0.3\leq x_{\text{F}}\leq 0.93$, red) nuclear modification factor in $p+A$ collisions at $\sqrt{s_{NN}}=38.7$ GeV, measured by the E866 collaboration [40, 80].
@@
  • At LHCb energies, $R_{pA}^{J/\psi}\sim 0.6$ at forward rapidity ($42.5<y<$ 4) [100, 103], stronger than predicted by nPDF shadowing. This suppression is consistent with observations by the ALICE collaboration [94, 97], which reported a similar trend in $J/\psi$ suppression at forward rapidity in $p+\mathrm{Pb}$ collisions at the same beam energy. In addition, the observed hierarchy of nuclear modification factors, $R_{pA}^{\psi{\rm(2S)}}<R_{pA}^{J/\psi}<R_{pA}^{\Upsilon(1S)}$ [110, 112], suggests the presence of nuclear effects beyond nPDFs alone.
@@
    Figure 4: The nuclear dependence of (top) Drell-Yan and (bottom) $J/\psi$ production as a function of $x_{2}$. The Drell-Yan data, $R_{pA}^{\rm DY}$, are from E772 [39], E866 [40], and E906 [44]. The $J/\psi$ data, presented as a function of the exponent $\alpha$, are from NA3 [73], E866 [80], PHENIX [153], ALICE [154], and LHCb [155]. Adapted from [42].
All these observations, across a range of $x_{F}$, $x_{2}$ and $p_{T}$, suggest the need to go beyond the conventional nPDF framework and demonstrate the energy dependence of CNM effects. We attempt to identify the physics mechanisms that could provide a complete and universal description of these data. The following sections introduce the CNM effects that may underlie the observed experimental trends and aim to address the three key questions outlined in Section I. We also propose a set of observables designed to disentangle contributions to the observed behavior from individual nuclear effects, thereby facilitating a more precise characterization of the nuclear modifications.
@@ III Disentangling nPDFs from saturation
If this suppression is genuinely nonperturbative, one must then rely on phenomenological models and possibly lattice methods, which currently face severe limitations. One such model is the Glauber-Gribov model [159, 160, 161, 162, 163] of (virtual) photon-nucleus total cross sections (structure functions) where the exchanged degrees of freedom are nonperturbative color singlet states called “pomerons” [164, 165] (or soft pomerons, to be distinguished from hard or BFKL pomerons [166]) which carry the quantum numbers of vacuum. In this approach, suppression of the nuclear cross section is due to the destructive interference between the phases of multiple pomeron exchanges. The nuclear parton distributions are thus determined at an initial scale $Q_{0}^{2}$ and then evolved in $Q^{2}$ via the DGLAP equations [167, 168]; this procedure is known as leading-twist shadowing. The evolution of the parton distribution functions is identical for nucleons and nuclei, with all nuclear effects (shadowing) encoded in the initial distributions. These nuclear PDFs, assumed to be universal, are then used within the collinear factorization framework to compute particle production cross sections in nuclear collisions.
@@
On the other hand, if weakly coupled QCD dynamics underlie the observed suppression, it should be calculable directly in terms of quark and gluon degrees of freedom. One such mechanism is gluon saturation, formulated as the Color Glass Condensate (CGC) effective field theory of small $x$ [169, 170, 171]. Such a saturated state of gluons, characterized by a relatively hard scale, the saturation scale, $Q_{s}\gg\Lambda_{\rm QCD}$, can be reached in a large nucleus already at moderately small values of $x$ ($<4-5\times 10^{-2}$) without significant evolution in $x$ whereas, in the proton, saturation is expected to be reached at much smaller values of $x$. In this regime of high gluon densities, $Q^{2}\lesssim Q_{s}^{2}$, nonlinear gluon recombination effects play a crucial role in taming the rapid growth of the gluon distribution at low $x$, leading to saturation of the distribution, as required by unitarity.
@@
The saturation scale, which delineates the transition between the dilute and dense regimes of QCD matter, depends on the atomic mass number $A$, the gluon momentum fraction $x$ (which coincides with Bjorken $x$ in structure functions but not necessarily in particle production), and the impact parameter of the collisions with the target nucleus or proton. Importantly, as an effective theory of QCD at small $x$, the CGC framework applies to a broad class of observables beyond structure functions [172, 173].
@@
Note that CGC dynamics can lead to both higher-twist and leading-twist suppression of nuclear cross sections. In the kinematic regime where small-$x$ evolution is significant but the momentum or virtuality remains well above the saturation scale ($Q^{2}\gg Q_{s}^{2}$), the cross section can be expanded in powers of $1/Q^{2}$, with the leading contribution corresponding to leading-twist dynamics. Although the system is dilute in this regime, leading-twist shadowing arises due to the presence of a saturation boundary separating the dense and dilute regimes. The extent of the suppression is sensitive to the dynamics near the saturation scale, which can be predicted within the CGC effective theory. Lastly, there has been significant progress toward calculating next-to-leading order (NLO) corrections for various observables using the CGC formalism which is crucial for precise predictions [174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209].